OpenStudy (anonymous):

Marla begins walking at 3 mi/h towards the library . Her friend meets her at the halfway point and drives her the rest of the way tot the library. The distance to the library is 4miles . How Many Miles did marla walk ? * Have Too Make An Equation * * Also Have Too Make A Sketch ( I Have The Sketch )*

OpenStudy (radar):

This problem has a lot of distracting data. Here is what is important in order to solve. The distance to the library is 4 miles. She (Maria) walked half the distance. As her friend met her at the halfway point.

OpenStudy (anonymous):

how would i put that in equation form

OpenStudy (omnomnom):

4/2= ? IDK IM HORRIBLE AT THIS

OpenStudy (radar):

Let x = distance Maria walked 2x=4 x=4/2 x = 2 miles

OpenStudy (omnomnom):

well i was close but radar is rightt ^_^

OpenStudy (radar):

You were correct @omnomnom

OpenStudy (omnomnom):

OMG REALLY? O.O

OpenStudy (radar):

I just put it in more of an "equation" like form. Hopefully that is what the instructor is looking for.

OpenStudy (omnomnom):

Oh... well im happy now ^_^ thank you :D

OpenStudy (anonymous):

thanks you guys really helped

OpenStudy (omnomnom):

:D welcome

OpenStudy (radar):

Yes, and good luck with your studies.